# All-in-one simulation-based inference

### v1

@@ Abstract
Figure 1: Capabilities of the Simformer: It can perform inference for simulators with a finite number of parameters or function-valued parameters (first column), it can exploit dependency structures of the simulator to improve accuracy (second column), it can perform inference for unstructured and missing data and for observation intervals (third column), and it provides an ‘all-in-one’ inference method that can sample all conditionals of the joint distribution, including posterior and likelihood (fourth column).
@@ 1 Introduction
Many methods for amortized SBI have been developed recently (Papamakarios & Murray, 2016; Lueckmann et al., 2017; Le et al., 2017; Greenberg et al., 2019; Papamakarios et al., 2019; Radev et al., 2020; Hermans et al., 2020; Glöckler et al., 2022; Deistler et al., 2022a; Simons et al., 2023). While these methods have different strengths and weaknesses, most of them also share limitations. First, they rely on structured, tabular data (typically ${\bm{\theta}},{\bm{x}}$ vectors). Yet, real-world datasets are often more messy (Shukla & Marlin, 2021): Irregularly sampled time series naturally arise in domains like ecology, climate science, and health and missing values often occur in real-world observations and are not easily handled by existing approaches. Second, the inputs of a simulator can correspond to a function of time or space, i.e., $\infty$-dimensional parameters (Chen et al., 2020; Ramesh et al., 2022). Existing amortized methods typically necessitate discretization, thereby limiting their applicability to a specific, often dense grid and precludes the evaluation of the parameter posterior beyond this grid. Third, they require specification of a fixed approximation task: the neural network can either target the likelihood (neural likelihood estimation, NLE, Papamakarios et al. (2019)) or the posterior (neural posterior estimation, NPE, Papamakarios & Murray (2016)). In practice, users might want to interactively explore both conditional distributions, investigate posteriors conditioned on subsets of data and parameters, or even explore different prior configurations. Fourth, while neural-network based SBI approaches are more efficient than classical ABC-methods (Lueckmann et al., 2021), they are still simulation-hungry. In part, this is because they target black-box simulators, i.e., do not require any access to the model’s inner workings. However, in practice, one has at least partial knowledge (or assumptions) about the structure of the simulator (i.e., its conditional independencies), but common SBI methods cannot exploit such knowledge. These limitations have prevented the application of SBI in interactive applications, in which properties of the task need to be changed on the fly.
@@ 3.1 A Tokenizer for SBI
The Simformer uses learnable vector embeddings for identifiers and condition states, as proposed in Weilbach et al. (2023). In cases where parameters or data are functions of space or time, the node identifier will comprise a shared embedding vector and a random Fourier embedding of the elements in the index set. Finally, specialized embedding networks, commonly used in SBI algorithms and trained end-to-end (Lueckmann et al., 2017; Chan et al., 2018; Radev et al., 2020), can be efficiently integrated here by condensing complex data into a single token. This reduces computational complexity but loses direct control over dependencies and condition states for individual data elements.
@@ 3.3 Simformer training and sampling
After having trained the Simformer, it can straightforwardly sample arbitrary conditionals (Fig. 3). We draw samples from the noise distribution and run the reverse diffusion process on all unobserved variables, while keeping observed variables constant at their conditioning value (Weilbach et al., 2023). Having access to all conditional distributions also allows us to combine scores and thereby perform inference for simulators with i.i.d. datapoints (Geffner et al., 2023). Similarly, we can use this fact to adapt to other prior or likelihood configurations post-hoc (see Appendix Sec. A1.4).
@@ 4.1 Benchmark tasks
Figure 4: Simformer performance on benchmark tasks. (a) Classifier Two-Sample Test (C2ST) accuracy between Simformer- and ground-truth posteriors. (b) C2ST between arbitrary Simformer-conditional distributions and their ground truth.
We evaluated performance in approximating posterior distributions across four benchmark tasks (Lueckmann et al., 2021). For each task, samples for ten ground-truth posteriors are available (Appendix Sec. A2.2), and we assessed performance as classifier two-sample test (C2ST) accuracy to these samples. Here, a score of 0.5 signifies perfect alignment with the ground truth posterior, and 1.0 indicates that a classifier can completely distinguish between the approximation and the ground truth. All results are obtained using the Variance Exploding SDE (VESDE); additional results using the Variance Preserving SDE (VPSDE) can be found in Appendix Sec. A3. See Appendix Sec. A2 for details on the exact parameterization.
@@
Next, we evaluated the ability of the Simformer to evaluate arbitrary conditionals. Arbitrary parameter and data conditions often vastly differ from the form of the posterior distribution, leading to a challenging inference task (Fig. 3). We performed inference on two of the benchmark tasks and established two new tasks with particularly interesting dependencies (Tree and HMM, details in Appendix Sec. A2.2). For each of the tasks, we generated ground truth posterior samples with Markov-Chain Monte-Carlo on 100 randomly selected conditional or full joint distributions. We found that, despite the complexity of these tasks, Simformer was able to accurately model all conditionals across all tasks (Fig. 4b). We note that, surprisingly, training solely on the posterior mask does not enhance performance relative to learning all conditional distributions (Appendix Sec. A3).
@@ 4.2 Lotka-Volterra: Inference with unstructured observations
Many measurements in science are made in an unstructured way. For example, measurements of the populations of prey and predator species in ecology might not always be made at the same time points, and even the number of observations that were made might differ between species. To demonstrate that Simformer can deal with such ‘unstructured’ datasets, we applied the method to the Lotka-Volterra model of ecology (Lotka, 1925; Volterra, 1926). The Lotka-Volterra model is a classic representation of predator-prey dynamics and is characterized by four global parameters, which govern the growth, hunting, and death rates of prey and predator. These populations evolve over time, guided by a set of coupled ordinary differential equations with Gaussian observation noise (details in Sec. A2.2). We note that, unlike Lueckmann et al. (2021), we perform inference for the full time-series and do not rely on summary statistics.
We trained Simformer on $10^{5}$ simulations and, after training, generated several synthetic observations. The first of these observations contained four measurements of the prey population, placed irregularly in time (green crosses in Fig. 5a). Simformer inferred the posterior distribution given these data. We found that the ground truth parameter set was indeed within regions of high posterior probability, and the Simformer posterior closely matched the ground truth posterior generated with MCMC(Fig. 5c, Appendix Sec. A2.2). We then used the ability of Simformer to sample from arbitrary
@@
conditional distribution to simultaneously generate posterior and posterior predictive samples without additional runs of the simulator. The posterior predictives of Simformer capture data and uncertainty in a realistic manner (Fig. 5a).
@@ 4.3 SIRD-model: Inference in infinite dimensional parameters
Figure 7: Inference in the Hodgkin-Huxley model. (a) Model schematic, observed voltage trace, and associated energy consumption. (b) Marginals of inferred posterior for four parameters. (c) Posterior predictive energy consumption from Simformer (blue) and from simulator outputs (green). (d) Posterior predictive samples from the simulator. (e) Marginals of inferred energy constrained posterior for four parameters. (f) Posterior predictive energy consumption from Simformer (blue) and from simulator outputs (green). Energy constraint as red line. (g) Posterior predictive samples from posterior in (e).
We generated synthetic measurements from infected, recovered, and deceased individuals at irregularly spaced time points and applied the Simformer to estimate the posterior distribution of parameters. The Simformer estimated realistic death and recovery rates and successfully recovers a time-dependent contact rate that aligns with ground truth observations (Fig. 6a). Indeed, as measurements of infections tend towards zero (around timestamp 25, Fig. 6a, orange), the Simformer-posterior for the contact rate increases its uncertainty. This is expected, as we cannot obtain conclusive insights about the contact rate in scenarios with negligible infections. Additionally, as we had also demonstrated on the Lotka-Volterra task, the ability of the Simformer to sample any conditional distribution allows to generate posterior predictive samples without running the simulator. These samples closely match the observed data, further demonstrating the accuracy of the Simformer.
Next, we demonstrate that the Simformer can accurately sample parameter-conditioned posterior distributions. We generated a synthetic observation consisting of four measurements of the time-dependent contact rate parameter and a single measurement of infected people. The resulting Simformer-posterior closely aligns with the parameter measurements, and its posterior predictives are aligned with the data.
Overall, these results demonstrate that the Simformer can tackle function-valued parameter spaces and perform accurate inference in these spaces and that its ability to sample arbitrary conditionals allows to incorporate parameter measurements into the inference procedure.
@@ 4.4 Hodgkin-Huxley model: Inference with observation intervals
Finally, we demonstrate that the Simformer can perform inference in a highly nonlinear model and that it can constrain the parameters to observation intervals with guided diffusion. For example, in neuroscience, it is desirable to obtain parameter configurations conditioned to experimental voltage measurements but also restricted by constraints such as lowering the metabolic cost below a particular threshold. Such additional constraints can be formalized as an observation intervals.
We will demonstrate the ability of Simformer to perform such inferences in an example from Neuroscience, the Hodgkin-Huxley simulator (Hodgkin & Huxley, 1952). This simulator describes the time course of action potentials along the membrane of neurons (Fig. 7a). The simulator has $7$ parameters and generates a noisy time series, which we reduced to summary statistics as in previous work (Gonçalves et al., 2020). In addition, we also record the metabolic cost consumed by the circuit and add it as an additional statistic (Appendix Sec. A2.2).
We first inferred the posterior distribution given only the summary statistics of the voltage (not the energy) with the Simformer, and we found that, consistent with prior work (Gonçalves et al., 2020), the posterior distribution has wide marginals for some parameters and narrow marginals for others (Fig. 7b). We then used Simformer’s ability to sample arbitrary conditionals and generate posterior predictives for energy consumption. The posterior predictive distribution of Simformer closely matched the posterior predictive distribution obtained by running the simulator, and the energy cost of different posterior samples varied significantly (Deistler et al., 2022b).
To identify energy-efficient parameter sets, we then defined an observation interval for the energy consumption (energy must be within the lowest 10% quantile of posterior predictives), and we used Simformer with guided diffusion to infer the posterior given voltage summary statistics and this constraint on energy consumption. The additional constraint on energy consumption significantly constrained the parameters posterior, in particular the maximal sodium and potassium conductances (Fig. 7e). We generated posterior predictive samples from this new posterior (via Simformer and by running the simulation) and found that their energy consumption indeed lies below the desired threshold (Fig. 7f, further details and results on guidance in Appendix Sec. A3.2, Fig. A7) Overall, Simformer can successfully recover the posterior distribution of highly nonlinear simulators. Simformer can condition on exact observations but also, by using guided diffusion, on nearly arbitrary constraints (see Appendix Fig. A3, Fig A8).
@@ 5 Discussion
We developed the Simformer, a new method for simulation-based amortized inference. The Simformer outperforms previous state-of-the-art methods (NPE) for posterior inference and simultaneously estimates all other conditionals. On tasks with notable independent structures, Simformer can be (on average across tasks and observations), one order of magnitude more simulation-efficient if equipped with a proper attention mask. The Simformer is significantly more flexible than previous out-of-the box inference frameworks and allows us to perform inference in $\infty$-dimensional parameter spaces, on unstructured and missing data. The Simformer makes it possible to sample arbitrary (or specified) conditional distributions of the joint distribution of parameters and data, including posterior and likelihood, thereby providing an ‘all-in-one’ inference method. These conditional distributions can be used to perform inference with parameter conditionals, or to obtain posterior predictive samples without running the simulator. Using diffusion guidance, one can also condition on intervals, which, e.g., can also be used to modify the prior without the need for retraining. Overall, the Simformer is an accurate and highly flexible inference method that opens up new possibilities for amortized inference methods in science and engineering.
@@
In this work, we focus on estimating all conditionals, a task that, within our framework, is roughly as complex as learning the joint distribution. In problems with a few parameters but high dimensional data (i.e. images or long time series), estimating the joint might be harder than just the posterior. In such cases, Simformer can simply be queried to target specific conditionals of interest (e.g. posterior and missing data posteriors).
@@ Conclusion
We developed the Simformer, a new method for amortized simulation-based inference. On benchmark tasks, it performs at least as well as existing methods that only target the posterior, although the Simformer estimates all conditional distributions. The Simformer is highly flexible and can jointly tackle multiple amortized inference tasks more effectively than previous methods.
@@ Acknowledgements
For an undirected, connected graph, all variables can depend on each other (given $l$ is large enough). This is a core argument by Weilbach et al. (2023) that an undirected graphical representation, given enough layers, is enough to faithfully represent all dependencies for any condition. Yet, this also diminishes any chance of correctly enforcing correct independencies beyond separating disconnected components. On the other hand, a directed acyclic graph will stay directed and acyclic. This property disallows modeling arbitrary dependencies, and this is why we have to dynamically adapt the mask to faithfully represent dependencies for arbitrary conditionals. We use the algorithm as proposed by Webb et al. (2018), which returns a minimal amount of edges we have to add to the directed graph to faithfully represent present dependencies (under certain topological ordering constraints). This is shown in Figure A1b. As expected for modeling the likelihood, no additional edges have to be introduced. On the other hand, to model the posterior distribution, we have to insert additional edges into the upper right corner. Note that this mask is sufficient to represent dependencies with a 1-layer transformer and thus adds too many edges in general. For Gaussian linear tasks, where $M_{E}$ stands as an idempotent matrix (i.e. $M_{E}^{2}=M_{E}$), resulting in $D=M_{E}$, this implies that all conditional independencies are correctly enforced, elucidating the substantial enhancement in accuracy. Even if dependencies are not exactly enforced, as observed by both our work and Weilbach et al. (2023), structured masks can enhance performance and computational complexity, particularly in the presence of notable independence structures. It is important to note that these dependencies are what is enforced by the model, not what is necessarily learned.
@@
$\nabla_{\theta_{t}}\log p_{t}(\theta_{t}|\mathbf{x}_{t})=\nabla_{\theta_{t}}\log p_{t}(\theta_{t})+\nabla_{\theta_{t}}\log p_{t}(\mathbf{x}_{t}|\theta_{t}).$ (1)
@@
$\nabla_{\theta_{t}}\log p_{t}(\mathbf{x}_{t}|\theta_{t})\approx s_{\phi}(\theta_{t},t|\mathbf{x}_{t})-s_{\phi}(\theta_{t},t),$ (2)
@@
$\nabla_{\theta_{t}}\log p_{t}^{\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}}(\theta_{t}|\mathbf{x}_{t})\approx\underbrace{\alpha_{1}\cdot(s_{\phi}(\theta_{t},t)+\beta_{1})}_{\text{Prior change}}+\underbrace{\alpha_{2}\cdot(s_{\phi}(\theta_{t},t|\mathbf{x}_{t})-s_{\phi}(\theta_{t},t)+\beta_{2})}_{\text{Likelihood change}}.$ (3)
@@
$\nabla_{z}\log\mathcal{N}(z;\mu_{0},\sigma_{0}^{2})=\frac{z-\mu_{0}}{\sigma_{0}^{2}},$
@@
In all our experiments, we sampled the condition mask $M_{C}$ as follows: At every training batch, we selected uniformly at random a mask corresponding to the joint, the posterior, the likelihood or two random masks. The random masks were drawn from a Bernoulli distribution with $p=0.3$ and $p=7$. In our experiments, we found this to work better than just random sampling and sufficiently diverse to still represent all the conditionals. The edge mask $M_{E}$ is chosen to match the generative process. The undirected variant was obtained by symmetrization.
@@
The prior for the parameter $\theta$ is a uniform distribution, denoted as $\mathcal{N}(0,0.1\cdot I)$. The data $x$ given $\theta$ is generated as per a Gaussian distribution $N(x|m_{\theta}=\theta,S=0.1\cdot I)$ with dimensionality of both $\theta$ and $x$ is $\mathbb{R}^{10}$.
@@
This task, commonly referenced in Approximate Bayesian Computation (ABC) literature (Sisson et al., 2007; Beaumont et al., 2009), involves inferring the common mean of a mixture of two-dimensional Gaussian distributions with distinct covariances. The task is defined as follows. The prior for the parameters $\theta$ is a uniform distribution, denoted as $U(-10,10)$. The data $x$ given $\theta$ is modeled as a mixture of two Gaussian distributions:
$x|\theta\sim 0.5\cdot\mathcal{N}(x|m_{\theta}=\theta,S=I)+0.5\cdot\mathcal{N}(x|m_{\theta}=\theta,S=0.01\cdot I)$
The parameter space $\theta$ and the data space $x$ are both in $\mathbb{R}^{2}$.

### v2

@@ Abstract
Figure 1: Capabilities of the Simformer: It can perform inference for simulators with a finite number of parameters or function-valued parameters (first column), it can exploit dependency structures of the simulator to improve accuracy (second column), it can perform inference for unstructured or missing data, for observation intervals (third column), and it provides an ‘all-in-one’ inference method that can sample all conditionals of the joint distribution, including posterior and likelihood (fourth column).
@@ 1 Introduction
Many methods for amortized SBI have been developed recently (Papamakarios & Murray, 2016; Lueckmann et al., 2017; Le et al., 2017; Greenberg et al., 2019; Papamakarios et al., 2019; Radev et al., 2020; Hermans et al., 2020; Glöckler et al., 2022; Boelts et al., 2022; Deistler et al., 2022a; Simons et al., 2023). While these methods have different strengths and weaknesses, most of them also share limitations. First, they often rely on structured, tabular data (typically ${\bm{\theta}},{\bm{x}}$ vectors). Yet, real-world datasets are often more messy (Shukla & Marlin, 2021): Irregularly sampled time series naturally arise in domains like ecology, climate science, and health sciences. Missing values often occur in real-world observations and are not easily handled by existing approaches. Second, the inputs of a simulator can correspond to a function of time or space, i.e., $\infty$-dimensional parameters (Chen et al., 2020; Ramesh et al., 2022). Existing amortized methods typically necessitate discretization, thereby limiting their applicability to a specific, often dense grid and precludes the evaluation of the parameter posterior beyond this grid. Third, they require specification of a fixed approximation task: the neural network can either target the likelihood (neural likelihood estimation, NLE, Papamakarios et al. (2019)) or the posterior (neural posterior estimation, NPE, Papamakarios & Murray (2016)). In practice, users might want to interactively explore both conditional distributions, investigate posteriors conditioned on subsets of data and parameters, or even explore different prior configurations. Fourth, while neural-network based SBI approaches are more efficient than classical ABC-methods (Lueckmann et al., 2021), they are still simulation-hungry. In part, this is because they target blackbox simulators, i.e., they do not require any access to the model’s inner workings. However, in practice, one has at least partial knowledge (or assumptions) about the structure of the simulator (i.e., its conditional independencies), but common SBI methods cannot exploit such knowledge. These limitations have prevented the application of SBI in interactive applications, in which properties of the task need to be changed on the fly.
@@ 3.1 A Tokenizer for SBI
The Simformer uses learnable vector embeddings for identifiers and condition states, as proposed in Weilbach et al. (2023). In cases where parameters or data are functions of space or time, the node identifier will comprise a shared embedding vector and a random Fourier embedding of the elements in the index set. Finally, specialized embedding networks, commonly used in SBI algorithms and trained end-to-end (Lueckmann et al., 2017; Chan et al., 2018; Radev et al., 2020), can be efficiently integrated here by condensing complex data into a single token (e.g. we demonstrate this on a gravitational waves example in Appendix Sec. A3.2). This reduces computational complexity but loses direct control over dependencies and condition states for individual data elements.
@@ 3.3 Simformer training and sampling
After having trained the Simformer, it can straightforwardly sample arbitrary conditionals (Fig. 3). We draw samples from the noise distribution and run the reverse diffusion process on all unobserved variables, while keeping observed variables constant at their conditioning value (Weilbach et al., 2023). Having access to all conditional distributions also allows us to combine scores and thereby perform inference for simulators with i.i.d. datapoints (Geffner et al., 2023). Similarly, we can use other score transformations to adapt to other prior or likelihood configurations post-hoc (see Appendix Sec. A1.4).
@@ 4.1 Benchmark tasks
Figure 4: Simformer performance on benchmark tasks. The suffices ”undirected graph” and ”directed graph” denote Simformer variants with structured attention based on the respective graphical models. (a) Classifier Two-Sample Test (C2ST) accuracy between Simformer- and ground-truth posteriors. (b) C2ST between arbitrary Simformer-conditional distributions and their ground truth.
We evaluated performance in approximating posterior distributions across four benchmark tasks (Lueckmann et al., 2021). For each task, samples for ten ground-truth posteriors are available (Appendix Sec. A2.2), and we assessed performance as classifier two-sample test (C2ST) accuracy to these samples. Here, a score of 0.5 signifies perfect alignment with the ground truth posterior, and 1.0 indicates that a classifier can completely distinguish between the approximation and the ground truth. All results are obtained using the Variance Exploding SDE (VESDE); additional results using the Variance Preserving SDE (VPSDE) can be found in Appendix Sec. A3. See Appendix Sec. A2 for details on the parameterization.
@@
Next, we evaluated the ability of the Simformer to evaluate arbitrary conditionals. Arbitrary parameter and data conditions often vastly differ from the form of the posterior distribution, leading to a challenging inference task (Fig. 3). We performed inference on two of the benchmark tasks and established two new tasks with particularly interesting dependencies (Tree and HMM, details in Appendix Sec. A2.2). For each of the tasks, we generated ground truth posterior samples with Markov-Chain Monte-Carlo on 100 randomly selected conditional or full joint distributions. We found that, despite the complexity of these tasks, Simformer was able to accurately model all conditionals across all tasks (Fig. 4b). We note that training solely on the posterior mask does not enhance performance relative to learning all conditional distributions (Appendix Sec. A3). Further, Simformer is well calibrated (Appendix Fig. A9, Fig. A10, Fig. A11, Fig. A12) and, in most cases, also superior with respect to the loglikelihood (Appendix Fig. A8).
@@ 4.2 Lotka-Volterra: Inference with unstructured observations
Many measurements in science are made in an unstructured way. For example, measurements of the populations of prey and predator species in ecology might not always be made at the same time points, and even the number of observations that were made might differ between species. To demonstrate that Simformer can deal with such ‘unstructured’ datasets, we applied the method to the ecological Lotka-Volterra model (Lotka, 1925; Volterra, 1926). The Lotka-Volterra model is a classic representation of predator-prey dynamics and is characterized by four global parameters, which govern the growth, hunting, and death rates of prey and predator. These populations evolve over time, guided by a set of coupled ordinary differential equations with Gaussian observation noise (details in Sec. A2.2).
@@
We note that, unlike Lueckmann et al. (2021), we perform inference for the full time-series and do not rely on summary statistics.
We trained Simformer on $10^{5}$ simulations and, after training, generated several synthetic observations. The first of these observations contained four measurements of the prey population, placed irregularly in time (green crosses in Fig. 5a).
Using Simformer, we inferred the posterior distribution given this data. We found that the ground truth parameter set was indeed within regions of high posterior probability, and the Simformer posterior closely matched the ground truth posterior generated with MCMC (Fig. 5c, Appendix Sec. A2.2). We then used the ability of Simformer to sample from arbitrary conditional distribution to simultaneously generate posterior and posterior predictive samples without additional runs of the simulator. The posterior predictives of Simformer capture data and uncertainty in a realistic manner (Fig. 5a).
@@ 4.3 SIRD-model: Inference in infinite dimensional parameters
Figure 7: Inference in the Hodgkin-Huxley model. (a) Model schematic, observed voltage trace, and associated energy consumption. (b) Marginals of inferred posterior for four parameters. (c) Posterior predictive energy consumption from Simformer (blue) and from simulator outputs (green). (d) Posterior predictive samples from the posterior in (c) using the simulator. (e) Marginals of inferred energy constrained posterior for four parameters. (f) Posterior predictive energy consumption from Simformer (blue) and from simulator outputs (green). Energy constraint as red line. (g) Posterior predictive samples from posterior in (e) using the simulator.
We generated synthetic measurements from infected, recovered, and deceased individuals at irregularly spaced time points and applied the Simformer to estimate the posterior distribution of parameters. The Simformer estimated realistic death and recovery rates and successfully recovers a time-dependent contact rate that aligns with ground truth observations (Fig. 6a). Indeed, as measurements of infections tend towards zero (around timestamp 25, Fig. 6a, orange), the Simformer-posterior for the contact rate increases its uncertainty. This is expected, as we cannot obtain conclusive insights about the contact rate in scenarios with negligible infections. Additionally, as we already demonstrated on the Lotka-Volterra task, the ability of the Simformer to sample any conditional distribution allows us to generate posterior predictive samples without running the simulator. These samples closely match the observed data, further demonstrating the accuracy of the Simformer.
Next, we demonstrate that the Simformer can accurately sample parameter-conditioned posterior distributions (Fig. 6b). We generated a synthetic observation consisting of four measurements of the time-dependent contact rate parameter and a single measurement of infected people. The resulting Simformer-posterior closely aligns with the parameter measurements, and its posterior predictives are aligned with the data. We evaluate the performance quantitatively by computing the expected coverage, which verified that the conditional distributions estimated by Simformer are indeed well-calibrated (see Fig. A13).
Overall, these results demonstrate that the Simformer can tackle function-valued parameter spaces and that its ability to sample arbitrary conditionals allows the incorporation of parameter measurements or assumptions into the inference procedure.
@@ 4.4 Hodgkin-Huxley model: Inference with observation intervals
Finally, we demonstrate that the Simformer can perform inference in a highly nonlinear model and that it can constrain the parameters to observation intervals with guided diffusion. For example, in neuroscience, it is desirable to obtain parameter configurations conditioned to experimental voltage measurements but also restricted by constraints such as lowering the metabolic cost (energy) below a particular threshold. Such additional constraints can be formalized as observation intervals.
We demonstrate the ability of Simformer to perform such inferences in an example from neuroscience, the Hodgkin-Huxley simulator (Hodgkin & Huxley, 1952). This simulator describes the time course of voltage along the membrane of neurons (Fig. 7a). The simulator has $7$ parameters and generates a noisy time series, which we reduced to summary statistics as in previous work (Gonçalves et al., 2020). In addition, we also record the metabolic cost consumed by the circuit and add it as an additional statistic (Appendix Sec. A2.2).
We first inferred the posterior distribution given only the summary statistics of the voltage (not the energy) with the Simformer, and we found that, consistent with prior work (Gonçalves et al., 2020), the posterior distribution has wide marginals for some parameters and narrow marginals for others (Fig. 7b). We then used Simformer’s ability to sample arbitrary conditionals and generate posterior predictives for energy consumption (Fig. 7c). The posterior predictive distribution of Simformer closely matched the posterior predictive distribution obtained by running the simulator (Fig. 7cd), and the energy cost of different posterior samples varied significantly (Deistler et al., 2022b).
To identify energy-efficient parameter sets, we then defined an observation interval for the energy consumption (energy must be within the lowest 10% quantile of posterior predictives), and we used Simformer with guided diffusion to infer the posterior given voltage summary statistics and this constraint on energy consumption. The additional constraint on energy consumption significantly constrained the parameters posterior, in particular the maximal sodium and potassium conductances (Fig. 7e). We generated posterior predictive samples from this new posterior (via Simformer and by running the simulation) and found that their energy consumption indeed lies below the desired threshold (Fig. 7f). Furthermore, the corresponding predictive voltage trace is still in agreement with observations (Fig. 7g). Additional details and results on guidance are in Appendix Sec. A3.3 (e.g. Fig. A15 for benchmarks on the accuracy of guidance).
Overall, Simformer can successfully recover the posterior distribution of highly nonlinear simulators. Simformer can condition on exact observations but also, using guided diffusion, on nearly arbitrary constraints (see Appendix Fig. A3, Fig A16).
@@ 5 Discussion
We developed the Simformer, a new method for simulation-based amortized inference. The Simformer outperforms previous state-of-the-art methods (NPE) for posterior inference and simultaneously estimates all other conditionals. On tasks with notable independent structures, Simformer can be (on average across tasks and observations), one order of magnitude more simulation-efficient if equipped with a proper attention mask. The Simformer is significantly more flexible than previous out-of-the box inference frameworks and allows us to perform inference in $\infty$-dimensional parameter spaces, on unstructured and missing data. The Simformer makes it possible to sample arbitrary (or specified) conditional distributions of the joint distribution of parameters and data, including posterior and likelihood, thereby providing an ‘all-in-one’ inference method. These conditional distributions can be used to perform inference with parameter conditionals, or to obtain posterior predictive samples without running the simulator. Using diffusion guidance, one can also condition on intervals, which, e.g., can be used to modify the prior without the need for retraining. Overall, the Simformer is an accurate and highly flexible inference method that opens up new possibilities for amortized inference methods in science and engineering.
@@
In this work, we focus on estimating all conditionals, a task that, within our framework, is roughly as complex as learning the joint distribution. In problems with a few parameters but high dimensional data (i.e. images or long time series), estimating the joint might be harder than just the posterior. In such cases, Simformer can simply be queried to target specific conditionals of interest (e.g., posterior and missing data posteriors, see Appendix Sec. A3.2 for an example on gravitational waves).
@@ Conclusion
We developed the Simformer, a new method for amortized simulation-based inference. On benchmark tasks, it performs at least as well or better as existing methods that only target the posterior, although the Simformer estimates all conditional distributions. The Simformer is highly flexible and can jointly tackle multiple amortized inference tasks more effectively than previous methods.
@@ Acknowledgements
For an undirected, connected graph, all variables can depend on each other (given $l$ is large enough). This is a core argument by Weilbach et al. (2023) that an undirected graphical representation, given enough layers, is enough to faithfully represent all dependencies for any condition. Yet, this also diminishes any chance of correctly enforcing correct independencies beyond separating disconnected components. On the other hand, a directed acyclic graph will stay directed and acyclic. This property disallows modeling arbitrary dependencies, and this is why we have to dynamically adapt the mask to faithfully represent dependencies for arbitrary conditionals. We use the algorithm as proposed by Webb et al. (2018), which returns a minimal amount of edges we have to add to the directed graph to faithfully represent present dependencies (under certain topological ordering constraints). This is shown in Figure A1b. As expected for modeling the likelihood, no additional edges have to be introduced. On the other hand, to model the posterior distribution, we have to insert additional edges into the upper right corner. Note that this mask is sufficient to represent dependencies with a 1-layer transformer and thus adds too many edges in general. For Gaussian linear tasks, where $M_{E}$ stands as an idempotent matrix (i.e. $M_{E}^{2}=M_{E}$), resulting in $D=M_{E}$, this implies that all conditional independencies are correctly enforced, explaining the substantial enhancement in accuracy. Even if dependencies are not exactly enforced, as observed by both our work and Weilbach et al. (2023), structured masks can enhance performance and computational complexity, particularly in the presence of notable independence structures. It is important to note that these dependencies are what is enforced by the model, not what is necessarily learned.
@@
$\nabla_{{\bm{\theta}}_{t}}\log p_{t}({\bm{\theta}}_{t}|\mathbf{x}_{t})=\nabla_{{\bm{\theta}}_{t}}\log p_{t}({\bm{\theta}}_{t})+\nabla_{{\bm{\theta}}_{t}}\log p_{t}(\mathbf{x}_{t}|\theta_{t}).$ (1)
@@
$\nabla_{{\bm{\theta}}_{t}}\log p_{t}(\mathbf{x}_{t}|{\bm{\theta}}_{t})\approx s_{\phi}({\bm{\theta}}_{t},t|\mathbf{x}_{t})-s_{\phi}({\bm{\theta}}_{t},t),$ (2)
@@
$\nabla_{{\bm{\theta}}_{t}}\log p_{t}^{\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}}({\bm{\theta}}_{t}|\mathbf{x}_{t})\approx\underbrace{\alpha_{1}\cdot(s_{\phi}({\bm{\theta}}_{t},t)+\beta_{1})}_{\text{Prior change}}+\underbrace{\alpha_{2}\cdot(s_{\phi}({\bm{\theta}}_{t},t|\mathbf{x}_{t})-s_{\phi}({\bm{\theta}}_{t},t)+\beta_{2})}_{\text{Likelihood change}}.$ (3)
@@
$\nabla_{x}\log\mathcal{N}(x;\mu_{0},\sigma_{0}^{2})=\frac{x-\mu_{0}}{\sigma_{0}^{2}},$
@@
In all our experiments, we sampled the condition mask $M_{C}$ as follows: At every training batch, we selected uniformly at random a mask corresponding to the joint, the posterior, the likelihood or two random masks. The random masks were drawn from a Bernoulli distribution with $p=0.3$ and $p=7$. In our experiments, we found this to work slightly better than just random sampling and sufficiently diverse to still represent all the conditionals. The edge mask $M_{E}$ is chosen to match the generative process (see Fig. A4). The undirected variant was obtained by symmetrization. Note that this is the only input we provide; additional necessary dependencies, e.g., due to conditioning, are algorithmically determined (see Sec. A1.1).
@@
The prior for the parameter ${\bm{\theta}}$ is a normal distribution $\mathcal{N}(0,0.1\cdot\mathbf{I})$. The data ${\bm{x}}$ given ${\bm{\theta}}$ is generated by a Gaussian distribution $\mathcal{N}({\bm{x}};{\bm{\theta}},0.1\cdot\mathbf{I})$. Both ${\bm{\theta}},{\bm{x}}\in\mathbb{R}^{10}$.
@@
This task, commonly referenced in Approximate Bayesian Computation (ABC) literature (Sisson et al., 2007; Beaumont et al., 2009), involves inferring the common mean of a mixture of two-dimensional Gaussian distributions with distinct covariances. The task is defined as follows. The prior for the parameters ${\bm{\theta}}$ is a uniform distribution, denoted as $\mathcal{U}(-10,10)$. The data ${\bm{x}}$ given ${\bm{\theta}}$ is modeled as a mixture of two Gaussian distributions:
${\bm{x}}|{\bm{\theta}}\sim 0.5\cdot\mathcal{N}({\bm{x}};{\bm{\theta}},\mathbf{I})+0.5\cdot\mathcal{N}(x;{\bm{\theta}},0.01\cdot\mathbf{I})$
The parameter space ${\bm{\theta}}$ and the data space ${\bm{x}}$ are both in $\mathbb{R}^{2}$.
Figure A4: Directed base masks for each of the tasks.The Lotka Volterra mask dynamically adapts to different input times, here just for randomly selected times.